OpenStudy (anonymous):

Please help I fan and give medals! If f(x) varies directly with x2, and f(x) = 96 when x = 4, find the value of f(2). A.192 B.72 C.32 D.24

OpenStudy (anonymous):

@ParthKohli AHHH! I NEED YOUR SMARTNESS!

Parth (parthkohli):

Again, we have an equation f(x) = c * x^2

OpenStudy (anonymous):

Right can you walk me through it?

Parth (parthkohli):

Sorry... OK, so first let's find c.

Parth (parthkohli):

We can make out

Parth (parthkohli):

c from the given data above

OpenStudy (anonymous):

f(2)=c times x^2

Parth (parthkohli):

So for x = 4, we have f(x) = 96. Let's plug that in the original equation: 96 = c * (4)^2

Parth (parthkohli):

Solve for c.

OpenStudy (anonymous):

um how do I do that I probably sound realy dumb right now. :(

Parth (parthkohli):

What is 4^2?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

then what?

Parth (parthkohli):

96 = c * 16

OpenStudy (anonymous):

right

Parth (parthkohli):

c = 6

OpenStudy (anonymous):

how?

Parth (parthkohli):

96 = c * 16 Divide both sides by 16

OpenStudy (anonymous):

ok now what do I do?

Parth (parthkohli):

So you finally have f(x) = 6x^2. Use it to get f(2)

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

I'm sorry I'm just really confused on this whole thing. :(

Parth (parthkohli):

I'm sorry. It's a pain to type on my phone. I promise i can do better on the laptop.

OpenStudy (anonymous):

oh its ok

Parth (parthkohli):

f(2) = 6 * (2)^2

OpenStudy (anonymous):

f(2)=24?

Parth (parthkohli):

Yus!

OpenStudy (anonymous):

ok so then thats my answer?

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

Yes

OpenStudy (anonymous):

ok thanks again can I bother you one more time to check one of my answers?